# SKA-Bench: A Fine-Grained Benchmark for Evaluating Structured Knowledge Understanding of LLMs

### v1

@@ Abstract
Although large language models (LLMs) have made significant progress in understanding Structured Knowledge (SK) like KG and Table, existing evaluations for SK understanding are non-rigorous (i.e., lacking evaluations of specific capabilities) and focus on a single type of SK. Therefore, we aim to propose a more comprehensive and rigorous structured knowledge understanding benchmark to diagnose the shortcomings of LLMs. In this paper, we introduce SKA-Bench, a Structured Knowledge Augmented QA Benchmark that encompasses four widely used structured knowledge forms: KG, Table, KG+Text, and Table+Text. We utilize a three-stage pipeline to construct SKA-Bench instances, which includes a question, an answer, positive knowledge units, and noisy knowledge units. To evaluate the SK understanding capabilities of LLMs in a fine-grained manner, we expand the instances into four fundamental ability testbeds: Noise Robustness, Order Insensitivity, Information Integration, and Negative Rejection. Empirical evaluations on 8 representative LLMs, including the advanced DeepSeek-R1, indicate that existing LLMs still face significant challenges in understanding structured knowledge, and their performance is influenced by factors such as the amount of noise, the order of knowledge units, and hallucination phenomenon. Our dataset and code are available at https://github.com/Lza12a/SKA-Bench.
@@ 1 Introduction
dels (LLMs) OpenAI (2023); Dubey et al. (2024), Structured Knowledge (SK), such as knowledge graphs (KG) Bollacker et al. (2008) and tables, still remain essential due to their systematic and rigorous organizational formats. On the one hand, structured knowledge is usually present in various real-world scenarios (e.g., financial reports with numerous tables Chen et al. (2021) and product knowledge graphs Wu et al. (2024)), thus serving as a significant knowledge base for existing LLM systems Liang et al. (2024); Wang et al. (2025). On the other hand, due to their well-organized structure and intensive knowledge characteristics, structured knowledge is also widely utilized to improve the inference-time performances of LLMs Li et al. (2024a, b); Guan et al. (2024). Consequently, evaluating the ability of LLMs to understand structured knowledge is a crucial research topic.
@@
Unlike common unstructured text understanding tasks Guo et al. (2023), LLMs still face significant challenges Fang et al. (2024) in understanding structured knowledge. This is because LLMs need to capture long-distance contextual dependencies as well as complex relationships and hierarchical structures from the given structured knowledge. However, existing benchmarks Pasupat and Liang (2015); Wu et al. (2025); Talmor and Berant (2018); Wu et al. (2024) for evaluating structured knowledge understanding suffer from limitations, including the lack of detailed reasoning path annotations or sufficiently long structured knowledge bases, making it difficult to thoroughly diagnose the shortcomings of LLMs in structured knowledge understanding. Moreover, these datasets primarily focus on single data types, including tables Pasupat and Liang (2015); Wu et al. (2025), knowledge graphs Talmor and Berant (2018), or hybrid Chen et al. (2020b); Wu et al. (2024) formats, which restrict their coverage and fail to fully reflect the comprehensive understanding abilities of the models. Therefore, there is an urgent need for a diverse and fine-grained dataset to comprehensively evaluate LLMs and identify potential bottlenecks in their structured knowledge understanding capabilities.
@@ 3.2.2 Iterative Positive Units Annotation
After each round of annotation, we query the LLM (utilizing DeepSeek-v3 DeepSeek-AI et al. (2024)) to determine whether annotated positive units can derive the answer to the given question. If the response is “No”, re-annotation is performed. The iterative annotation process continues until more than 95% of the samples receive a “Yes” response, at which point the iteration is terminated.
@@ 3.2.3 Noisy Units Construction
First, we leverage LLMs with existing SKA-QA instances to generate noisy units. To ensure the diversity of synthesized units, we alternately use GPT-4o OpenAI (2023) and DeepSeek-v3 DeepSeek-AI et al. (2024) during this process. Meanwhile, we also need to ensure that the synthesized noisy units do not affect the correctness of the answers. To achieve this, we prompt LLM (utilizing DeepSeek-v3) with QA and positive units to derive the “conditions” that must be satisfied by the rows for answering the question. LLM then verifies whether the generated noisy units meet these “conditions”. If the response is “Yes”, the noisy units need to be re-generated by LLMs. After the noise synthesis process, three human experts conduct a manual review of Table subset and Table+Text subset to evaluate whether the synthetic noise is unsafe and affect the original answers. The review results show that the accuracy rate is 92.5%, and erroneous noise has been deleted.
@@ 3.4 Testbeds Construction
[…]
@@
[…]
@@ 4.1 Experimental Settings
Our evaluation is based on popular large language models (LLMs) with a context window of at least 24k tokens. Our evaluated LLMs include advanced general-purpose LLMs: DeepSeek-v3 DeepSeek-AI et al. (2024), GPT-4o OpenAI (2023), DeepSeek-R1 DeepSeek-AI et al. (2025) and common open-source LLMs: Llama-3.1-8B-Instruct Dubey et al. (2024), Qwen2.5-7B-Instruct Team (2024), GLM4-9B-Chat Zeng et al. (2024), and Mistral-7B-Instruct-v0.3 Jiang et al. (2023a). Moreover, we also evaluate the table-specific open-source LLM TableGPT-2 Su et al. (2024), which are trained based on Qwen2.5-7B.
@@ 5 Conclusion
Although SKA-Bench offers a more comprehensive and rigorous benchmark for evaluating structured knowledge understanding of LLMs, certain limitations warrant careful consideration, as summarized below. (1) SKA-Bench is limited to English only and does not yet capture the performances of LLMs in understanding structured knowledge across multiple languages. (2) Constrained by resource limitations, although our SKA-Bench instances have the capability to construct longer structured knowledge bases (even >64k tokens), we have not yet explored the performance of LLMs at this scale.
@@
In this paper, we construct SKA-Bench, which is expanded and modified based on the existing 6 structured knowledge understanding evaluation datasets. Moreover, we incorporate manual annotation and manual synthetic data verification to ensure that it does not violate any ethics.
@@ Appendix A Original Datasets Details
Appendix B Dataset Construction Details
@@
Appendix C Evaluation Prompt Template

### v2

@@ Abstract
Although large language models (LLMs) have made significant progress in understanding Structured Knowledge (SK) like KG and Table, existing evaluations for SK understanding are non-rigorous (i.e., lacking evaluations of specific capabilities) and focus on a single type of SK. Therefore, we aim to propose a more comprehensive and rigorous structured knowledge understanding benchmark to diagnose the shortcomings of LLMs. In this paper, we introduce SKA-Bench, a Structured Knowledge Augmented QA Benchmark that encompasses four widely used structured knowledge forms: KG, Table, KG+Text, and Table+Text. We utilize a three-stage pipeline to construct SKA-Bench instances, which includes a question, an answer, positive knowledge units, and noisy knowledge units. To evaluate the SK understanding capabilities of LLMs in a fine-grained manner, we expand the instances into four fundamental ability testbeds: Noise Robustness, Order Insensitivity, Information Integration, and Negative Rejection. Empirical evaluations on 8 representative LLMs, including the advanced DeepSeek-R1, indicate that existing LLMs still face significant challenges in understanding structured knowledge, and their performance is influenced by factors such as the amount of noise, the order of knowledge units, and hallucination phenomenon. Our dataset and code are available at https://github.com/zjukg/SKA-Bench.
@@ 1 Introduction
dels (LLMs) OpenAI (2023); AI@Meta (2024), Structured Knowledge (SK), such as knowledge graphs (KG) Bollacker et al. (2008) and tables, still remain essential due to their systematic and rigorous organizational formats. On the one hand, structured knowledge is usually present in various real-world scenarios (e.g., financial reports with numerous tables Chen et al. (2021) and product knowledge graphs Wu et al. (2024)), thus serving as a significant knowledge base for existing LLM systems Liang et al. (2024); Wang et al. (2025). On the other hand, due to their well-organized structure and intensive knowledge characteristics, structured knowledge is also widely utilized to improve the inference-time performances of LLMs Li et al. (2024a, b); Guan et al. (2024). Consequently, evaluating the ability of LLMs to understand structured knowledge is a crucial research topic.
@@
Unlike common unstructured text understanding tasks Guo et al. (2023), LLMs still face significant challenges Fang et al. (2024); Liu et al. (2025) in understanding structured knowledge. This is because LLMs need to capture long-distance contextual dependencies as well as complex relationships and hierarchical structures from the given structured knowledge. However, existing benchmarks Pasupat and Liang (2015); Wu et al. (2025); Talmor and Berant (2018); Wu et al. (2024) for evaluating structured knowledge understanding suffer from limitations, including the lack of detailed reasoning path annotations or sufficiently long structured knowledge bases, making it difficult to thoroughly diagnose the shortcomings of LLMs in structured knowledge understanding. Moreover, these datasets primarily focus on single data types, including tables Pasupat and Liang (2015); Wu et al. (2025), knowledge graphs Talmor and Berant (2018), or hybrid Chen et al. (2020b); Wu et al. (2024) formats, which restrict their coverage and fail to fully reflect the comprehensive understanding abilities of the models. Therefore, there is an urgent need for a diverse and fine-grained dataset to comprehensively evaluate LLMs and identify potential bottlenecks in their structured knowledge understanding capabilities.
@@ 3.2.2 Iterative Positive Units Annotation
After each round of annotation, we query the LLM (utilizing DeepSeek-v3 DeepSeek-AI (2024)) to determine whether annotated positive units can derive the answer to the given question. If the response is “No”, re-annotation is performed. The iterative annotation process continues until more than 95% of the samples receive a “Yes” response, at which point the iteration is terminated.
@@ 3.2.3 Noisy Units Construction
First, we leverage LLMs with existing SKA-QA instances to generate noisy units. To ensure the diversity of synthesized units, we alternately use GPT-4o OpenAI (2023) and DeepSeek-v3 DeepSeek-AI (2024) during this process. Meanwhile, we also need to ensure that the synthesized noisy units do not affect the correctness of the answers. To achieve this, we prompt LLM (utilizing DeepSeek-v3) with QA and positive units to derive the “conditions” that must be satisfied by the rows for answering the question. LLM then verifies whether the generated noisy units meet these “conditions”. If the response is “Yes”, the noisy units need to be re-generated by LLMs. After the noise synthesis process, three human experts conduct a manual review of Table subset and Table+Text subset to evaluate whether the synthetic noise is unsafe and affect the original answers. The review results show that the accuracy rate is 92.5%, and erroneous noise has been deleted.
@@ 3.4 Testbeds Construction
$\bullet$ Noise Robustness. Here, we define noise as the remaining triples in the KG subgraph or the irrelevant rows in the table. We incorporate noise units of varying proportions into the positive knowledge units as the knowledge base to evaluate whether the LLM can robustly provide accurate answers. Considering the differences in the token counts across different knowledge units, we use the total token length as the split standard to construct test sets. Specifically, we construct four test sets {1k, 4k, 12k, 24k} for the Table and KG subsets, and three test sets {4k, 12k, 24k} for the Table+Text and KG+Text subsets, with the detailed statistics shown in Table 2. Additionally, to eliminate the influence of the knowledge unit order, we randomly shuffle the SK units in the KG and text units with a random seed of 42, while preserving the original order of the SK units in the Table.
@@
$\bullet$ Order Insensitivity. SK representation naturally does not depend on any specific order. And in retrieval-augmented scenarios Fan et al. (2024), the order of retrieved knowledge units tends to be disrupted. Therefore, we expect LLMs to be order-insensitive when understanding SK and capturing the semantic relationships between SK units. In this testbed, we provide SK bases with different permutations of SK units to test whether the LLM is sensitive to order. For SK units in KG and textual units, we position the positive knowledge units at the beginning, randomized positions, and the end of the knowledge base, denoting them as {prefix, random, suffix}. For SK units in Table, we additionally introduce the original table order, denoted as {original, prefix, random, suffix}. Furthermore, we standardize the test sets to a scale of 4k tokens for Table and KG subsets, and 12k for Table+Text and KG+Text subsets.
$\bullet$ Information Integration. This ability requires LLMs to integrate multiple knowledge units to answer questions, including the integration of multiple SK units and the integration of heterogeneous data (SK+Text) units. Therefore, this testbed focuses on analyzing the performance of LLMs under these two settings. Specifically, we divide our dataset based on the number of knowledge units required to answer each question {2, 3, 4, more than 4} to evaluate the information integration capability of LLMs. Regarding dataset scale and order, we standardize the test set to a scale of 4k tokens for the Table and KG subsets, and 12k tokens for Table+Text and KG+Text subsets. Meanwhile, we randomly shuffle (with random seed 42) the SK units in KG and text units while preserving the original order of SK units in the Table subset.
@@ 4.1 Experimental Settings
Our evaluation is based on popular large language models (LLMs) with a context window of at least 24k tokens. Our evaluated LLMs include advanced general-purpose LLMs: DeepSeek-v3 DeepSeek-AI (2024), GPT-4o OpenAI (2023), DeepSeek-R1 DeepSeek-AI (2025) and common open-source LLMs: Llama-3.1-8B-Instruct AI@Meta (2024), Qwen2.5-7B-Instruct Team (2024), GLM4-9B-Chat Zeng et al. (2024), and Mistral-7B-Instruct-v0.3 Jiang et al. (2023a). Moreover, we also evaluate the table-specific open-source LLM TableGPT-2 Su et al. (2024), which are trained based on Qwen2.5-7B.
@@ 5 Conclusion
Although SKA-Bench offers a more comprehensive and rigorous benchmark for evaluating structured knowledge understanding of LLMs, certain limitations warrant careful consideration, as summarized below. (1) SKA-Bench is limited to English only and does not yet capture the performances of LLMs in understanding structured knowledge across multiple languages. (2) Constrained by resource limitations, although our SKA-Bench instances have the capability to construct longer structured knowledge bases (even >64k tokens), we have not yet explored the performance of LLMs at this scale. (3) Potential biases introduced by dataset selection are inevitable. We have made effort to minimize this impact through careful dataset selection and evaluation design. For dataset selection, SKA-Bench adopts widely used and well-recognized benchmark datasets for each respective data type (e.g., WebQSP, CWQ, WTQ, TableBench, StarK-Amazon, and HybridQA). These datasets focus on general domains and do not require extensive domain-specific knowledge, which helps to mitigate the bias introduced by domain dependence. For evaluation design, our primary focus is on whether LLMs can comprehensively understand various types of structured knowledge, rather than investigating the “performance differences between different modalities”. (4) Beyond its role as a structured knowledge understanding dataset, SKA-Bench can be effectively extended to various task scenarios, such as Text2Query (i.e., Text2SQL Guo et al. (2019) and Text2SPARQL Hu et al. (2018)), structured knowledge retrieval Mahalingam et al. (2024), and evaluations of knowledge-augmented systems Liang et al. (2024); Wang et al. (2025). These potential evaluation directions remain unexplored.
@@
In this paper, we construct SKA-Bench, which is expanded and modified based on the existing 6 structured knowledge understanding evaluation datasets. These datasets have stated that there are no ethical concerns. Moreover, we also incorporate manual annotation and manual synthetic data verification to ensure that it does not violate any ethics.
Acknowledgments
This work is founded by National Natural Science Foundation of China (NSFC62306276/NSFCU23B2055/NSFCU19B2027), Zhejiang Provincial Natural Science Foundation of China (No. LQ23F020017), Yongjiang Talent Introduction Programme (2022A-238-G), and Fundamental Research Funds for the Central Universities (226-2023-00138). This work was supported by Ant Group.
@@ Appendix A Original Datasets Details
Type 1-hop 2-hop 3-hop $\mathbf{\geq}$4-hop
KG subset 7.59% 22.07% 16.26% 54.09%
KG+Text subset 7.04% 10.91% 9.76% 72.29%
Table 6: Hop count distributions in KG subset and KG+Text subset.
Model KG KG+Text
Hard Noise Mixed Noise Easy Noise Std. Hard Noise Mixed Noise Easy Noise Std.
LLaMA3.1-8B 55.12 58.19 60.23 2.10 65.68 67.02 66.96 0.62
Qwen2.5-7B 58.36 60.00 67.82 4.13 74.96 76.51 77.15 0.92
DeepSeek-v3 71.09 73.93 74.28 1.43 77.01 77.12 77.65 0.28
GPT-4o 72.10 73.42 74.82 1.11 76.58 77.38 77.63 0.45
Table 7: Performance of four representative LLMs under different noise intensities on KG and KG+Text subsets.
Appendix B Effects of Noise Distribution in KG-related Subsets
Due to the structural complexity of knowledge graphs, the degree of noise can vary across different triples. We hypothesize that hop count is closely related to semantic similarity, and this relationship may affect the noise distribution, thereby impacting model performance. To investigate this, we conduct an in-depth analysis of the hop count distribution in the existing KG subsets and KG+Text subsets, and provide experimental results to support this hypothesis. As shown in Table 6, we present the detailed hop count distributions for the KG and KG+Text subsets. It is worth noting that we report the actual distributions in SKA-Bench rather than the original datasets, since some subgraphs were randomly truncated during the graph extraction process.
Thanks to the high flexibility of SKA-Bench, we further present results under different noise intensities to more thoroughly evaluate the noise robustness of LLMs. In our experimental setup, the context size is consistently controlled at 4k, and knowledge units from the knowledge graph are arranged in a random order. Noise is categorized as “Hard Noise” ($\leq$2 hops), “Mixed Noise” (same as in main text), and “Easy Noise” ($\geq$4 hops). We report the performance of four representative LLMs, as shown in Table 7. The experimental results show that the variation in model performance is significantly correlated with the intensity of noise. Comparatively, stronger LLMs demonstrate greater robustness to noise, and this conclusion also holds when the noise level is expanded horizontally (as in the noise robustness testbed of the main text).
Appendix C Dataset Construction Details
@@
Appendix D Evaluation Prompt Template